given that n is a positive even integer, 5n + 4 will always be divisible

A. 5n

B.5

C. 2

D. 4


1
Expert's answer
2022-02-23T12:29:49-0500

Let "n=2k,k=1,2,3,..."

Then "5n+4=10k+4=2(5k+2)=2m, m=7,8,9,..."


Therefore "5n + 4" will always be divisible by "2."

C. 2